OpenStudy (4hxlols):

Given tanØ =4/3 and sin(-4/5) find cosØ A. 1/3 B. - 1/3 C. 3/5 D. - 3/5 E. -√3/5

OpenStudy (antiprime):

if cosØ = -3/5 and sinØ < 0 find tanØ (nearest fourth decimal digit) ------- If cos and sin are both negative theta is in QIII. ---- Since cos = x/r = -3/5, x = -3 and r = 5 ---- Solve for "y": y = sqrt[r^2 - x^2] = sqrt[25-9] = sqrt(16) = 4 --- But y is negative in QIII, so y = -4 ----- Tan is positive in QIII: tan(theta) = y/x = -4/-3 = 4/3

OpenStudy (antiprime):

thats a dif question but im showing u how to work a problem like that

OpenStudy (4hxlols):

could you give an example where the value of sin isn't 0 so i can compare them easier

OpenStudy (4hxlols):

i follow you up until "solve for y"

OpenStudy (freckles):

question seems a little off

OpenStudy (freckles):

Did you mean: "Given tan(Ø)=4/3 and sin(Ø)=-4/5 find cos(Ø)"

OpenStudy (4hxlols):

here u go freckles

OpenStudy (freckles):

wow they made a huge type-o then

OpenStudy (4hxlols):

ive found so many errors in the basic math on plato it worrys me when i get to the harder stuff like this, especially when others see the mistakes too :(

OpenStudy (freckles):

I bet you they meant to write "Given tan(Ø)=4/3 and sin(Ø)=-4/5 find cos(Ø)"

OpenStudy (freckles):

it shouldn't say sin(-4/5) this tells us nothing

OpenStudy (4hxlols):

well in that case can you help me solve that problem and ill put the answer to that in?

OpenStudy (freckles):

\[\tan(\theta)=\frac{\sin(\theta)}{\cos(\theta)}\] use this

OpenStudy (freckles):

replace tan(theta) with 4/3 and replace sin(theta) with -4/5 and solve for cos(theta)

OpenStudy (freckles):

if you want you can replace cos(theta) with u if that makes things easier to see and helps you solve

OpenStudy (4hxlols):

soooooo -3/5?

OpenStudy (freckles):

\[\frac{4}{3}=\frac{\frac{-4}{5}}{u} \\ \frac{4}{3}=\frac{-4}{5 u} \\ \\ \text{ cross multiply } \\ 20u=-12 \\ u=\frac{-12}{20}=\frac{-3}{5}\] yep

OpenStudy (freckles):

but seriously I hope you see why the sin(-4/5) is a type-o

OpenStudy (4hxlols):

thank you so much, and i do thats why the first answer seemed confusing. its sad that plato isnt formatted correctly though :/

OpenStudy (freckles):

very sad

OpenStudy (freckles):

you would think tons of people would have looked it over

OpenStudy (freckles):

and corrected the software

OpenStudy (4hxlols):

especially when we are expected to only use the knowledge given to us by plato, asking for help like i did is considered "cheating" lol